The tires of a 15700 N automobile are inflated to a pressure of 172000 Pa (25 psi). Assuming that each tire supports 1⁄4 of the weight of the car, calculate the area of contact between a tire and the road.
1
Expert's answer
2020-03-26T12:17:25-0400

We have our pressure in each of the four tires that withstands car's weight:


"4p(\text{1 wheel})=\frac{F}{4A(\text{1 wheel})},\\
\space\\
A(\text{1 wheel})=\frac{F}{16p(\text{1 wheel})}\\
\space\\=\frac{15700}{16\cdot172000}=5.7\cdot10^{-3}\text{ m}^2,"

or 57 square centimeters. This is a rectangle 21.5 cm wide - average wheel width and 2.7 cm long - the flat part that you see on the round surface of the wheel:



Source of the image: https://www.autotrader.ca/newsfeatures/20180507/under-inflated-tires-are-dangerous-and-cost-you-money/